Numerical Analogues of Aronson’s Sequence

Benoit Cloitre 13 rue Pinaigrier Tours 37000, FRANCE (Email: abmt@wanadoo.fr)

N. J. A. Sloane AT&T Shannon Labs Florham Park, NJ 07932–0971, USA (Email: njas@research.att.com)

Matthew J. Vandermast 53 Piaget Avenue Clifton, NJ 07011–1216, USA (Email: ghodges14@msn.com)

Abstract

Aronson’s sequence 1, 4, 11, 16, … is defined by the English sentence “t is the first, fourth, eleventh, sixteenth, … letter of this sentence.” This paper introduces some numerical analogues, such as: a(n) is taken to be the smallest positive integer greater than a(n − 1) which is consistent with the condition “n is a member of the sequence if and only if a(n) is odd.” This sequence can also be characterized by its “square”, the sequence a^{(2)}(n) = a(a(n)), which equals 2n + 3 for n ≥ 1. There are many generalizations of this sequence, some of which are new, while others throw new light on previously known sequences.

1. Introduction

Aronson’s sequence is defined by the English sentence “t is the first, fourth, eleventh, sixteenth, … letter of this sentence (not counting spaces or commas),” and is a classic example of a self-referential sequence ([3], [8], sequence M3406 in [13], A5224 in [12]). It is somewhat unsatisfactory because of the ambiguity in the English names for numbers over 100 — for example, some people say “one hundred and one”, while others say “one hundred one.” Another well-known example is Golomb’s sequence, in which the n^{th} term G(n) (for n ≥ 1) is the number of times n appears in the sequence (A1462 in [12]):

1, 2, 2, 3, 3, 4, 4, 4, 5, 5, 5, 6, 6, 6, 6, 7, 7, 7, 7, 8, …

There is a simple formula for G(n): it is the nearest integer to (and approaches)

ϕ^{2−ϕ}n^{ϕ−1} ,

where ϕ = (1 + \sqrt{5})/2 ([5], [6, Section E25]).

Additional examples can be found in Hofstadter’s books [7], [8] and in [6] and [12]. However, the sequence {a(n)} mentioned in the Abstract appears to be new, as do many of the other sequences we will discuss. We will also give new properties of some sequences that have been studied elsewhere.

Section 2 discusses the sequence mentioned in the Abstract, and also introduces the “square” of a sequence. Some simple generalizations (non-monotonic, “even” and “lying” versions) are described in Section 3. The original sequence is based on examination of the sequence modulo 2. In Section 4 we consider various “mod y” generalizations. Section 5 extends both the original sequence and the “mod y” generalizations by defining the “Aronson transform” of a sequence. Finally, Section 6 briefly considers the case when the rule defining the sequence depends on more than one term.

There are in fact a large number of possible generalizations and we mention only some of them here. We have not even analyzed all the sequences that we do mention. In some cases we just list the first few terms and invite the reader to investigate them himself. We give the identification numbers of these sequences in [12] — the entries there will be updated as more information becomes available.

We have also investigated sequences arising when (2) is replaced by the following rule: s(1) = x, s(n) = s(n − 1) + y if n is already in the sequence, s(n) = s(n − 1) + z otherwise, for specified values of x, y, z. This work will be described elsewhere [4].

Notation. “Sequence” here usually means an infinite sequence of nonnegative numbers. “Monotonically increasing” means that each term is strictly greater than the previous term. ℙ = {1, 2, 3, …}, ℕ = {0, 1, 2, 3, …}.

2. n is in the sequence if and only if a(n) is odd

Let the sequence a(1), a(2), a(3), … be defined by the rule that a(n) is the smallest positive integer > a(n − 1) which is consistent with the condition that

“n is a member of the sequence if and only if a(n) is odd.”  (1)

The first term, a(1), could be 1, since 1 is odd and 1 would be in the sequence. It could also be 2, since then 1 would not be in the sequence (because the terms must increase) and 2 is even. But we must take the smallest possible value, so a(1) = 1. Now a(2) cannot be 2, because 2 is even. Nor can a(2) be 3, for then 2 would not be in the sequence but a(2) would be odd. However, a(2) = 4 is permissible, so we must take a(2) = 4, and then 2 and 3 are not in the sequence.

So a(3) must be even and > 4, and a(3) = 6 works. Now 4 is in the sequence, so a(4) must be odd, and a(4) = 7 works. Continuing in this way we find that the first few terms are as follows (this is A79000):

\begin{matrix}n : & 1 & 2 & 3 & 4 & 5 & 6 & 7 & 8 & 9 & 10 & 11 & 12 & ⋯ \\ a(n) : & 1 & 4 & 6 & 7 & 8 & 9 & 11 & 13 & 15 & 16 & 17 & 18 & ⋯ \end{matrix}

Once we are past a(2) there are no further complications, a(n − 1) is greater than n, and we can, and therefore must, take

a(n) = a(n − 1) + ϵ ,  (2)

where ϵ is 1 or 2 and is given by:

a(n − 1) even a(n − 1) odd
n in sequence  1  2
n not in sequence  2  1

The gap between successive terms for n ≥ 3 is either 1 or 2.

The analogy with Aronson’s sequence is clear. Just as Aronson’s sentence indicates exactly which of its terms are t’s, {a(n)} indicates exactly which of its terms are odd.

We proceed to analyze the behavior of this sequence.

First, all odd numbers ≥ 7 occur. For suppose 2t + 1 is missing. Therefore a(i) = 2t, a(i + 1) = 2t + 2 for some i ≥ 3. From the definition, this means i and i + 1 are missing, implying a gap of at least 3, a contradiction.

Table I shows the first 72 terms, with the even numbers underlined.

Examining the table, we see that there are three consecutive numbers, 6, 7, 8, which are necessarily followed by three consecutive odd numbers, a(6) = 9, a(7) = 11, a(8) = 13. Thus 9 is present, 10 is missing, 11 is present, 12 is missing, and 13 is present. Therefore the sequence continues with a(9) = 15 (odd), a(10) = 16 (even), …, a(13) = 19 (odd), a(14) = 20 (even). This behavior is repeated for ever. A run of consecutive numbers is immediately followed by a run of the same length of consecutive odd numbers.

Let us define the k^{th} segment (for k ≥ 0) to consist of the terms a(n) with n = 9 · 2^{k} − 3 + j where −3 · 2^{k} ≤ j ≤ 3 · 2^{k} − 1. In the table the segments are separated by vertical lines. The first half of each segment, the terms where j < 0, consists of consecutive numbers given by a(n) = 12 · 2^{k} − 3 + j; the second half, where j ≥ 0, consists of consecutive odd numbers given by a(n) = 12 · 2^{k} − 3 + 2j. We can combine these formulae, obtaining an explicit description for the sequence:

a(1) = 1, a(2) = 4,

and subsequent terms are given by

a(9 · 2^{k} − 3 + j) = 12 · 2^{k} − 3 + \frac{3}{2}j + \frac{1}{2}|j|  (3)

for k ≥ 0, −3 · 2^{k} ≤ j < 3 · 2^{k}.

The structure of this sequence is further revealed by examining the sequence of first differences, Δa(n) = a(n + 1) − a(n), n ≥ 1, which is

3, 2, 1, 1, 1, 2, 2, 2, 1^{6}, 2^{6}, 1^{12}, 2^{12}, 1^{24}, 2^{24}, …  (4)

n : 1 2 3 4 5 6 7 8 9 10
a(n) : 1 \underline{4} \underline{6} 7 \underline{8} 9 11 13 15 \underline{16}
n : 11 12 13 14 15 16 17 18 19 20
a(n) : 17 \underline{18} 19 \underline{20} 21 23 25 27 29 31
n : 21 22 23 24 25 26 27 28 29 30
a(n) : 33 \underline{34} 35 \underline{36} 37 \underline{38} 39 \underline{40} 41 \underline{42}
n : 31 32 33 34 35 36 37 38 39 40
a(n) : 43 \underline{44} 45 47 49 51 53 55 57 59
n : 41 42 43 44 45 46 47 48 49 50
a(n) : 61 63 65 67 69 \underline{70} 71 \underline{72} 73 \underline{74}
n : 51 52 53 54 55 56 57 58 59 60
a(n) : 75 \underline{76} 77 \underline{78} 79 \underline{80} 81 \underline{82} 83 \underline{84}
n : 61 62 63 64 65 66 67 68 69 70
a(n) : 85 \underline{86} 87 \underline{88} 89 \underline{90} 91 \underline{92} 93 95
n : 71 72 ⋯
a(n) : 97 99 ⋯

Table I: The first 72 terms of the sequence “n is in the sequence if and only if a(n) is odd.”

(A79948), where we have written 1^{m} to indicate a string of m 1’s, etc. The oscillations double in length at each step.

Segment 0 begins with an even number, 6, but all other segments begin with an odd number, 9 · 2^{k} − 3. All odd numbers occur in the sequence except 3 and 5. The even numbers that occur are 4, 6, 8 and all numbers 2m with

9 · 2^{k−1} − 1 ≤ m ≤ 6 · 2^{k} − 2, k ≥ 1 .

The sequence of differences, (4), can be constructed from the words in a certain formal language (cf. [11]). Let the alphabet be A = {1, 2, 3}, and let A^{∗} denote the set of strings of elements from A. We define a mapping θ from A^{∗} to A^{∗} by the rules θ(1) = 2, θ(2) = 1, 1 . Then (4) is the concatenation

S_{−1}, S_{0}, S_{1}, S_{2}, …,  (5)

where

S_{−1} = {3, 2}, S_{0} = {1, 1, 1}, S_{k+1} = θ(S_{k}) for k ≥ 0 .  (6)

To prove this, note that for n ≥ 3, a difference of 2 only occurs in {a(n)} between a pair of odd numbers. Suppose a(i) = 2j + 1, a(i + 1) = 2j + 3; then a(2j + 1) = 2x + 1 (say), a(2j + 2) = 2x + 2, a(2j + 3) = 2x + 3, producing two differences of 1. Similarly, if there is a difference of 1, say a(i) = j, a(i + 1) = j + 1, then a(j) = 2x + 1, a(j + 1) = 2x + 3, a difference of 2.

The ratio n/a(n), which is the fraction of positive integers in the sequence that are less than or equal to a(n), rises from close to 2/3 at the beginning of segment k (assuming k is large), reaches a maximum 3/4 at the midpoint of the segment, then falls back to 2/3 at the end of the segment. It is not difficult to show that if n is chosen at random in the k^{th} segment then the average value of the fraction of numbers in the sequence at that point approaches

\frac{3}{4} − \frac{1}{4} log \frac{32}{27} = 0.7075 …

for large k.

The sequence has an alternative characterization in terms of its “square.”

The square of a sequence s = {s(n) : n ≥ n_{0}} is given by s^{(2)} = {s(s(n)) : n ≥ n_{0}}. If s is monotonically increasing so is s^{(2)}.

Lemma 1. Let s be monotonically increasing. Then n (≥ n_{0}) is in the sequence s if and only if s(n) is in the sequence s^{(2)}.

Proof. If n is in the sequence, n = s(i) for some i ≥ n_{0}, and s(n) = s(s(i)) is in s^{(2)}. Conversely, if s(n) ∈ s^{(2)}, s(n) = s(s(i)) for some i ≥ n_{0}, and since s is monotonically increasing, n = s(i).

For our sequence a = {a(n)}, examination of Table I shows that a^{(2)} = {1, 5, 7, 9, 11, …} = {1} ∪ 2ℙ + 3. This can be used to characterize a. More precisely, the sequence can be defined by: a(1) = 1, a(2) = 4, a(3) = 6 and, for n ≥ 4, a(n) is the smallest positive integer which is consistent with the sequence being monotonically increasing and satisfying a(a(n)) = 2n + 3 for n ≥ 2.

This is easily checked. Once the first three terms are specified, the rule a(a(n)) = 2n + 3 determines the remaining terms uniquely.

In fact that rule also forces a(2) to be 4, but it does not determine a(3), since there is an earlier sequence {a′(n)} (in the lexicographic sense) satisfying a′(1) = 1, a′(a′(n)) = 2n + 3 for n ≥ 2, namely

1, 4, 5, 7, 9, 10, 11, 12, 13, 15, 17, 19, 21, 22, …,

(A80596), and given by a′(1) = 1,

a′(6 · 2^{k} − 3 + j) = 8 · 2^{k} − 3 + \frac{3}{2}j + \frac{1}{2}|j|  (7)

for k ≥ 0, −2^{k+1} ≤ j < 2^{k+1}.

As the above examples show, the square of a sequence does not in general determine the sequence uniquely. A better way to do this is provided by the “inverse Aronson transform”, discussed in Section 5.

3. First generalizations

The properties of {a(n)} given in Section 2 suggest many generalizations, some of which will be discussed in this and the following sections.

(3.1) Non-monotonic version. If we replace “a(n) > a(n − 1)” in the definition by “a(n) is not already in the sequence”, we obtain a completely different sequence, suggested by J. C. Lagarias [10]: b(n), n ≥ 1, is the smallest positive integer not already in the sequence which is consistent with the condition that “n is a member of the sequence if and only if b(n) is odd.” This sequence (A79313) begins:

\begin{matrix}n : & 1 & 2 & 3 & 4 & 5 & 6 & 7 & 8 & 9 & 10 \\ b(n) : & 1 & 3 & 5 & \underline{2} & 7 & \underline{8} & 9 & 11 & 13 & \underline{12} \\ n : & 11 & 12 & 13 & 14 & 15 & 16 & 17 & 18 & 19 & 20 \\ b(n) : & 15 & 17 & 19 & \underline{16} & 21 & 23 & 25 & \underline{20} & 27 & 29 \end{matrix}

The even members are underlined. The behavior is simpler than that of {a(n)}, and we leave it to the reader to show that, for n ≥ 5, b(n) is given by

b(4t − 2) = 4t ,
b(4t − 1) = 6t − 3 ,
b(4t) = 6t − 1 ,
b(4t + 1) = 6t + 1 .

All odd numbers occur. The only even numbers are 2 and 4t, t ≥ 2. (The square b^{(2)} is not so interesting.)

(3.2) “Even” version. If instead we change “odd” in the definition of {a(n)} to “even”, we obtain a sequence c which is best started at n = 0: c(n), n ≥ 0, is the smallest nonnegative integer > c(n − 1) which is consistent with the condition that

“n is a member of the sequence if and only if c(n) is even.”  (8)

This is A79253: 0, 3, 5, 6, 7, 8, 10, 12, 14, 15, . . .. It is easily seen that c(n) = a(n+1)−1 for n ≥ 0, so there is nothing essentially new here. Also c^{(2)} = {0} ∪ 2ℙ + 4.

(3.3) The “lying” version. The lying version of Aronson’s sequence is based on the completely false sentence “t is the second, third, fifth, sixth, seventh, … letter of this sentence.” The sentence specifies exactly those letters that are not t’s, and produces the sequence (A81023) 2, 3, 5, 6, 7, 8, 9, 10, 11, 12, . . ..

Just as {a(n)} is an analogue of Aronson’s sequence, we can define an analogue {d(n) : n ≥ 1} of this sequence by saying that: d(n) is the smallest positive integer > d(n − 1) such that the condition “n is in the sequence if and only if d(n) is odd” is false. Equivalently, the condition “either n is in the sequence and d(n) is even or n is not in the sequence and d(n) is odd” should be true. The resulting sequence (A80653) begins 2, 4, 5, 6, 8, 10, 11, 12, 13, 14, . . .. We will give an explicit formula for d(n) in the next section.

A related sequence is also of interest. Let {d′(n)} be defined by d′(1) = 2, and, for n > 1, d′(n) is the smallest integer greater than d′(n − 1) such that the condition “n and d′(d′(n)) have opposite parities” can always be satisfied. One can show that this is the sequence

2, 4, 5, 7, 8, 9, 11, 12, 13, 14, 16, …

(A14132), the complement of the triangular numbers (A217), with d′(n) = n + (nearest integer to \sqrt{2n}).

4. The “mod m” versions

Both {a(n)} and {c(n)} are defined modulo 2. Another family of generalizations is based on replacing 2 by some fixed integer y ≥ 2. To this end we define a sequence {s(n) : n ≥ n_{0}} by specifying a starting value s(n_{0}) = s_{0}, and the condition that n is in the sequence if and only if s(n) ≡ z (mod y), where y and z are given.

Although we will not digress to consider this here, it is also of interest to see what happens when “if and only if” in the definition is replaced by either “if” or “only if.” (We mention just one example. The above sequence {d(n)}, prefixed by d(0) = 0, can be defined as follows: d(n) is the smallest nonnegative number > d(n − 1) such that the condition “n (n ≥ 0) is in the sequence only if d(n) is even” is satisfied.)

We saw in the previous section that {a(n)} can also be characterized by the property that its square a^{(2)}(n) = a(a(n)) is equal to 2n + 3 for n ≥ 2 (together with some appropriate initial conditions). This too can be generalized by specifying that the sequence {s(n)} satisfy s(s(n)) = yn + z, for given values of y and z. The two generalizations are related, but usually lead to different sequences. The s(s(n)) family of generalizations will connect the present investigation with several sequences that have already appeared in the literature. There are too many possibilities for us to give a complete catalogue of all the sequences that can be obtained from these generalizations. Instead we will give a few key examples and one general theorem. Many other examples can be found in [12].

A simple “mod 3” generalization is: e(1) = 2, and, for n > 1, e(n) is the smallest integer > e(n − 1) which is consistent with the condition that

“n is a member of the sequence if and only if e(n) is a multiple of 3.”  (9)

This turns out to be James Propp’s sequence

2, 3, 6, 7, 8, 9, 12, 15, 18, 19, …,

which appeared as sequence M0747 in [13] (A3605 in [12]). Propp gave a different (although equivalent) definition involving the square of the sequence: {e(n)} is the unique monotonically increasing sequence satisfying e(e(n)) = 3n for all n ≥ 1. Michael Somos [14] observed that this sequence satisfies

e(3n) = 3e(n),
e(3n + 1) = 2e(n) + e(n + 1),
e(3n + 2) = e(n) + 2e(n + 1).  (10)

An analysis similar to that for {a(n)} leads to the following explicit formula, which appears to be new:

e(2 · 3^{k} + j) = 3^{k+1} + 2j + |j| ,  (11)

for k ≥ 0 and −3^{k} ≤ j < 3^{k}.

A sequence closely related to {e(n)} had earlier been studied by Arkin et al. [2, Eq. (12)]. This is the sequence e′(n) = e(n) − n (A6166). Arkin et al. give a recurrence similar to (10). The sequences defined by recurrences s(s(n)) = kn for k ≥ 4 have recently been studied by Allouche et al. [1].

The sequence {e(n)} can be generalized as follows.

Theorem 2. Let y and z be integers of opposite parity satisfying

y ≥ 2, z ≥ 2 − y .  (12)

Then there is a unique monotonically increasing sequence \{f(n)\} satisfying f(1) = \frac{1}{2}(y+ z + 1) and f(f(n)) = yn + z for n > 1. It is given by

f (c_{1}y^{k} − \frac{z}{y − 1} + j) = c_{2}y^{k+1} − \frac{z}{y − 1} + \frac{y + 1}{2}j + \frac{y − 1}{2}|j| ,  (13)

for k ≥ 0, where

−\frac{y + z − 1}{2}y^{k} ≤ j < \frac{y + z − 1}{2}y^{k}

and

c_{1} = \frac{(y + 1)(y + z − 1)}{2(y − 1)}, c_{2} = \frac{y + z − 1}{y − 1} .

Proof. It is easy to see that the sequence {s(n)} defined by

s(n) = \frac{(y + 1)n − 2f(n) + z}{y − 1}  (14)

is a “saw-tooth” sequence of the form

0, 1, 2, 3, …, 3, 2, 1, 0, 1, 2, 3, 4, 5, 6, …, 6, 5, 4, 3, 2, 1, 0, 1, 2, 3, …,

where the 0’s occur at

n = 1 + \frac{(y + z − 1)(y^{k} − 1)}{y − 1} for k ≥ 0 .

Consequently the k-th “tooth” 0, 1, 2, 3, …, 3, 2, 1, 0 contains (y + z − 1)y^{k} terms and reaches a maximal value of (y+z−1)y^{k}/2. Then from simple considerations of symmetry, for any k ≥ 0 and any j such that |j| ≤ (y + z − 1)y^{k}/2, we have

s (1 + \frac{(y + z − 1)(y^{k} − 1)}{y − 1} + \frac{(y + z − 1)y^{k}}{2} + j) = \frac{(y + z − 1)}{2}y^{k} − |j| .

In other words,

s (\frac{(y + z − 1)(y + 1)}{2(y − 1)}y^{k} − \frac{z}{y − 1} + j) = \frac{(y + z − 1)}{2}y^{k} − |j| .  (15)

Equation (13) follows from (14) and (15).

It can be shown (we omit the details) that this sequence can also be defined by: f(1) = (y + z + 1)/2, and, for n > 1, f(n) is the smallest integer > f(n − 1) which is consistent with the condition that “n is a member of the sequence if and only if f(n) belongs to the set

[2, …, \frac{1}{2}(y + z − 1)] ∪ \{iy + z : i ≥ 1\} .”  (16)

If (y + z − 1)/2 ≤ 1, the first set in (16) is to be omitted.

Examples. Setting y = 3, z = 0 in the theorem produces {e(n)}.

Setting y = 2, z = 1 yields another interesting “mod 2” sequence. This is the sequence {g(n) : n ≥ 1} that begins

2, 3, 5, 6, 7, 9, 11, 12, 13, 14, …

(A80637). It has the following properties:

(i) By definition, this is the unique monotonically increasing sequence {g(n)} satisfying g(1) = 2, g(g(n)) = 2n + 1 for n ≥ 2.

(ii) n is in the sequence if and only if g(n) is an odd number ≥ 3.

(iii) The sequence of first differences is (A79882):

1, 2, 1^{2}, 2^{2}, 1^{4}, 2^{4}, 1^{8}, 2^{8}, 1^{16}, 2^{16}, ….

(iv)

g(3 · 2^{k} − 1 + j) = 2 · 2^{k+1} − 1 + \frac{3}{2}j + \frac{1}{2}|j| ,

for k ≥ 0, −2^{k} ≤ j < 2^{k} (from (13)).

(v) g(2n) = g(n) + g(n − 1) + 1, g(2n + 1) = 2g(n) + 1, for n ≥ 1 (taking g(0) = 0). (vi) The original sequence {a(n)} satisfies a(3n) = 3g(n), a(3n + 1) = 2g(n) + g(n +

1), a(3n + 2) = g(n) + 2g(n + 1), for n ≥ 1.

(vii) The “lying version” of Section 3 is given by d(n) = g(n + 1) − 1 for n ≥ 1. (viii) Let g′(n) = g(n) + 1. The sequence {g′(n) : n ≥ 2} was apparently first

discovered by C. L. Mallows, and is sequence M2317 in [13] (A7378 in [12]). This is the unique monotonically increasing sequence satisfying g′(g′(n)) = 2n. An alternative description is: g′(n) (for n ≥ 2) is the smallest positive integer > g′(n − 1) which is consistent with the condition that

“n is a member of the sequence if and only if g′(n) is an even number ≥ 4”.

(17)

Note that, although (8) and (17) are similar, the resulting sequences {c(n)} and {g′(n)} are quite different. g′ is not directly covered by Theorem 2, and we admit that we have not been able to identify the largest family of sequences which can be described by formulae like (3), (7), (11), (13).

The sequence {h(n)} defined by: h(1) = 2, and, for n > 1, h(n) is the smallest positive integer > h(n − 1) which is consistent with the condition that “n is a member of the sequence if and only if h(n) is a multiple of 6”:

2, 6, 7, 8, 9, 12, 18, 24, 30, 31, …,

(A80780), shows that such simple rules do not hold in general. We can characterize the sequence of first differences in a manner similar to (5), (6): the alphabet is now A = {1, 2, …, 6}, and we define a mapping θ from A^{∗} to A^{∗} by the rules θ(i) = 1, 1, …, 1, 7−i (with i − 1 1’s followed by 7 − i), for i = 1, …, 6. Then the sequence of differences of {h(n)} is S_{0}, S_{1}, S_{2}, …, where S_{0} = {4}, S_{k+1} = θ(S_{k}) for k ≥ 0. However, it appears that no formula similar to (3) holds for h(n).

We end this “mod m” section with two interesting “mod 4” sequences. The even numbers satisfy s(s(n)) = 4n, and the odd numbers satisfy s(s(n)) = 4n + 3. But there are lexicographically earlier sequences with the same properties. The “pseudoeven numbers” {i(n) : n ≥ 0} are defined by the property that i(n) is the smallest nonnegative integer > i(n − 1) and satisfying i(i(n)) = 4n (A80588):

0, 2, 4, 5, 8, 12, 13, 14, 16, 17, …

We analyze this sequence by describing the sequence of first differences, which are

2, 2, 1, 3, 4, 1, 1, 2, 1, 1, 1, 1, 4, 4, 1, 3, …

After the initial 2, 2, 1, this breaks up into segments of the form

3 S_{k} 2 T_{k} ,

where T_{k} is the reversal of

1^{1} 4^{2} 1^{4} 4^{8} 1^{16} 4^{32} ⋯ 4^{2^{2k−1}}1^{2^{2k}}

and S_{k} is the reversal of

1^{2} 4^{1} 1^{8} 4^{4} 1^{32} 4^{16} ⋯ 1^{2^{2k−1}}4^{2^{2k−2}} .

The “pseudo-odd numbers”, i′(n), are similarly defined by i′(i′(n)) = 4n + 3:

1, 3, 4, 7, 11, 12, 13, 15, 16, 17, …

(A80591), and satisfy i′(n) = i(n + 1) − 1.

5. The Aronson transform

A far-reaching generalization of both the original sequence and the “mod m” extensions of the previous section is obtained if we replace “odd number” in the definition of {a(n)} by “member of β”, where β is some fixed sequence.

More precisely, let us fix a starting point n_{0}, which will normally be 0 or 1. Let β = {β(n) : n ≥ n_{0}} be an infinite monotonically increasing sequence of integers ≥ n_{0} with the property that its complement (the numbers ≥ n_{0} that are not in β) is also infinite. Then the sequence α = {α(n) : n ≥ n_{0}} given by: α(n) is the smallest positive integer > α(n − 1) which is consistent with the condition that

“n is in α if and only if α(n) is in β”

is called the Aronson transform of β.

Theorem 3. The Aronson transform exists and is unique.

Proof. For ease of discussion let us call the numbers in β “hot”, and those in its complement “cold.” We will specify the transform α, leaving to the reader the easy verification that this has the desired properties, in particular that there are no contradictions.

The proof is by induction. First we consider the initial term α(n_{0}). If n_{0} is hot, α(n_{0}) = n_{0}. If n_{0} is cold, α(n_{0}) = smallest cold number ≥ n_{0} + 1.

For the induction step, suppose α(n) = k for n > n_{0}.

Case (i), k = n. If n + 1 is hot then α(n + 1) = n + 1. If n + 1 is cold then α(n + 1) = smallest cold number ≥ n + 2.

Case (ii), k > n. If k = n + 1 then α(n + 1) = smallest hot number ≥ n + 2. If k > n + 1 then if n + 1 is hot, α(n + 1) = smallest hot number ≥ k + 1, while if n + 1 is cold, α(n + 1) = smallest cold number ≥ k + 1.

In certain cases it may be appropriate to specify some initial terms in α to get it started properly.

Examples. Of course taking β to be the odd numbers (with n_{0} = 1) leads to our original sequence {a(n)}, and the even numbers (with n_{0} = 0) lead to {c(n)} of Section 3.

If we take β to be the triangular numbers we get 1, 4, 5, 6, 10, 15, 16, 17, 18, 21, … (A79257); the squares give 1, 3, 4, 9, 10, 11, 12, 13, 16, 25, … (A79258); the primes give 4, 6, 8, 11, 12, 13, 14, 17, 18, 20, … (A79254); and the lower Wythoff sequence (A201), in which the n^{th} term is ⌊nϕ⌋, gives 1, 5, 7, 10, 11, 13, 14, 15, 18, 19, … (A80760).

Taking the Aronson transform of {a(n)} itself we get 1, 3, 4, 6, 10, 11, 12, 14, 22, 23, … (A79325). [12] contains several other examples.

The inverse transform may be defined in a similar way. Given an infinite monotonically increasing sequence α = {α(n) : n ≥ n_{0}} of numbers ≥ n_{0}, such that its complement (the numbers ≥ n_{0} that are not in α) is also infinite, its inverse Aronson transform is the sequence β = {β(n) : n ≥ n_{0}} such that the Aronson transform of β is α.

Theorem 4. The inverse Aronson transform exists and is unique.

Proof. We establish this by giving a simple algorithm to construct the inverse transform. We illustrate the algorithm in Table II by applying it to the sequence of squares, α = {n^{2} : n ≥ 0}.

Form a table with four rows. In the first row place the numbers n = n_{0}, n_{0} + 1, n_{0} + 2, …, and in the second row place the sequence α(n_{0}), α(n_{0} + 1), . . .. The third row contains what we will call the “hot” numbers: these will comprise the elements of the inverse transform. The fourth row are the “cold” numbers, which are the complement of the hot numbers.

The third and fourth rows are filled in as follows. If n is in (resp. not in) the sequence α, place α(n) in the n-th slot of the hot (resp. cold) row.

To complete the table we must fill in the empty slots. Suppose we are at column n, where we have placed α(n) in one of the two slots. Let l_{n} be the largest number mentioned in columns n_{0}, …, n − 1 in the hot or cold rows. Then we place the numbers l_{n} + 1, …, α(n) − 1 in the empty slot in column n, with the single exception that if n is not in the sequence and l_{n} = n − 1 then we place n in the cold slot rather than the hot slot. (This is illustrated by the position of 2 in the fourth row of Table II.) We leave it to the reader to verify that the entries in the “hot” row form the inverse Aronson transform β.

\begin{matrix}n & 1 & 2 & 3 & 4 & 5 & 6 & 7 & 8 & 9 \\ α_{n} & 1 & 4 & 9 & 16 & 25 & 26 & 49 & 64 & 81 \\ “hot” & 1 & 3 & 5–8 & 16 & 17–24 & 26–35 & 37–48 & 50–63 & 81 \\ “cold” & − & 2, 4 & 9 & 10–15 & 25 & 36 & 49 & 64 & 65–80 \end{matrix}

Table II: Computation of the inverse Aronson transform of the squares. The “hot” numbers comprise the transform.

It follows from Lemma 1 that β contains the members of α^{(2)}, but in general β ≠ α^{(2)}. The additional terms in β make it possible to recover α uniquely from β.

Examples. As shown in Table II, the inverse Aronson transform of the squares (A10906) is

1; 3; 5, 6, 7, 8; 16; 17, …, 24; 26, 27, ….

This consists of a number of segments (separated here by semicolons). For k ≥ 1 the k^{th} segment is {k^{2}} if k is a square, or {(k − 1)^{2} + 1, …, k^{2} − 1} if k is not a square, except that the second segment = {3}.

The inverse transform of the primes is 3, 5, 6, 11, 12, 17, 18, 20, 21, 22, … (A80759)

— this has a similar decomposition into segments.

The inverse transform of the lower Wythoff sequence is 1, 4, 6, 7, 9, 10, 12, 14, 15, 17, … (A80746). This consists of the numbers ⌊ϕk⌋+k−1 (k ≥ 1) and ⌊2ϕk⌋+k−1 (k ≥ 2). The inverse transform of our original sequence {a(n)} is the sequence of odd numbers (whereas, as we saw in Section 2, a^{(2)} omits 3).

In general (because of the above algorithm), the inverse Aronson transforms are easier to describe than the direct transforms.

6. More complicated conditions

Finally, we may make the condition for n to be in the sequence depend on the values of several consecutive terms a(n), a(n + 1), …, a(n + τ), for some fixed τ. To pursue this further would take us into the realm of one-dimensional cellular automata (cf. [9], [15]), and we will mention just two examples.

q(n) is the smallest positive integer > q(n − 1) which is consistent with the condition that “n is in the sequence if and only if q(n) is odd and q(n − 1) is even” (A79255):

1, 4, 6, 9, 12, 15, 18, 20, 23, 26, 28, …

The gaps between successive terms are always 2 or 3. Changing the condition to “… both q(n) and q(n + 1) are odd” gives A79259:

1, 5, 6, 10, 11, 15, 19, 20, 24, 25, …

Acknowledgements

We thank J. C. Lagarias for some helpful comments, and the referee for a very careful reading of the manuscript.

References

[1] J.-P. Allouche, N. Rampersad and J. Shallit, On integer sequences whose first iterates are linear, Preprint, 2003.

[2] J. Arkin, D. C. Arney, L. S. Dewald and W. E. Ebel, Jr., Families of recursive sequences, J. Recreational Math., 22 (No. 22, 1990), 85–94.

[3] J. K. Aronson, quoted by D. R. Hofstadter in Metamagical Themas [8], p. 44.

[4] B. Cloitre, N. J. A. Sloane and M. J. Vandermast, Variations on a sequence of Recamán, in preparation, 2003.

[5] S. W. Golomb, Problem 5407, Amer. Math. Monthly, 73 (1966), 674; 74 (1967), 740–743.

[6] R. K. Guy, Unsolved Problems in Number Theory, Springer-Verlag, NY, 2nd ed., 1994.

[7] D. R. Hofstadter, Gödel, Escher, Bach: An Eternal Golden Braid, Vintage Books, NY, 1980.

[8] D. R. Hofstadter, Metamagical Themas, Basic Books, NY, 1985.

[9] A. Ilachinski, Cellular Automata, World Scientific, River Edge, NJ, 2001.

[10] J. C. Lagarias, Personal communication, 2003.

[11] M. Lothaire, Combinatorics on Words, Addison–Wesley, Reading, MA, 1983.

[12] N. J. A. Sloane, The On-Line Encyclopedia of Integer Sequences, published electronically at www.research.att.com/∼njas/sequences/, 2003.

[13] N. J. A. Sloane and S. Plouffe, The Encyclopedia of Integer Sequences, Academic Press, 1995.

[14] M. Somos, Personal communication, 2000.

[15] S. Wolfram, A New Kind of Science, Wolfram Media, 2002.

2000 Mathematics Subject Classification: Primary 11B37. Keywords: self-describing sequence, Aronson sequence, square of sequence

(Concerned with sequences \underline{A005224}, \underline{A001462} \underline{A079000} \underline{A079948} \underline{A080596} \underline{A079313}

\underline{A079253} \underline{A081023} \underline{A080653} \underline{A079325} \underline{A079257} \underline{A079258} \underline{A079254} \underline{A014132} \underline{A000217} \underline{A003605}
\underline{A006166} \underline{A080637} \underline{A079882} \underline{A007378} \underline{A080780} \underline{A080588} \underline{A080591} \underline{A000201} \underline{A080760} \underline{A010906}

\underline{A080759} \underline{A080746} \underline{A079255} \underline{A079259}.)

Received March 31, 2003; revised version received July 2, 2003. Published in Journal of Integer Sequences, July 4, 2003.

Return to Journal of Integer Sequences home page.